When the line y = 2x + 3 is reflected in the line y = x + 1, what is the equation of the reflected line?
Answer #1

I to mirror (lmle LG y = 2x+3. (a,b) Tyy=x+1 (6,K) (63) is an- random point on Reflected line y = 2x+3 Intersection of y=xto